OpenStudy (anonymous):

pls help.. a long straight rope lying on a floor is pulled along its length with uniform acceleration .If the velocities of the front & back ends of the rope when it passes through a particular point on the floor are 2 m/s & 4 m/s respectively, the velocity of the midpoint of the rope when it passes through the same point is a. 3 b. sqrt.10

OpenStudy (anonymous):

If v is the velocity of the midpoint. $$2*a*\frac l2 =v^2-(2)^2=(4)^2-v^2$$ which gives $$v=\sqrt{10} ms^{-1}$$